Reports 1-1 of 1 Clear search Modify search
takafumi.ushiba - 19:47 Friday 20 August 2021 (17974) Print this report
PF OLDAMP filter was tested

Continued work from klog17965.

I engaged MN_PFOLDAMP_Y filter to damp 4.1 Hz mode of suspension and succeeded.
I also check the performance of MN_MNOLDAMP_{R,P,Y} filters with photosensor loops and they seems work correctly with PS loops.

After the Ikeda-san's update work of Type-A model (klog17969), I started to design the damping loop of PF OpLev signals.
Designed filter was implementd at FM1 of MN_MNPFOLDAMP_Y filter bank with the name of "damp_temp".
I also put the PLANT of the suspension without any damping at FM10 of the same filter bank.

Figure 1 shows the transfer function from MN_TEST_Y_EXC to MN_OPLEV_WIT_Y, PF_OPLEV_WIT_Y, TM_OPLEV_WIT_Y, and MN_DAMP_Y_IN.
The measurement conditions are no damping, with only PS damp, with MNOL damp and PS damp, and with MNOL damp and PFOL damp.
Both MNOL damp + PS damp and MNOL damp + PFOL DAMP can damp all modes below 10Hz well, while the remaining peak can be observed when using only PS loop.

After that, I checked the long-term stability of the condition both PS damp and MN OLDAMP were engaged at the same time for R, P, Y degrees of freedom.
it seems good so far (about 30 min), so I keep the current condition tonight and see what happens.

When I worked, I stopped guardian because I'm not sure whether FLOAT state works well or not for Type-A suspension but of course it is not good situation.
So, making FLOAT state for Type-A suspension is an urgent task in my opinion.

Images attached to this report
Comments to this report:
takafumi.ushiba - 11:43 Saturday 21 August 2021 (17976) Print this report

I checked the status of lock from yesterday.
Even though there are many gliches in the feedback signals, the lock itself continues for more than half days.

So, I suggested that engaging MNOLDAMP_{R,P,Y} with PS damp as a default damping configuration at ALIGNED state.

Images attached to this comment
Search Help